# Branch length statistics in phylogenetic trees under constant-rate birth-death models

### v2

@@ VIII Branch length distributions conditioned on the population size at the time of observation
First we go backwards in time from the time of observation to the time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation is described by a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{n,N}^{\mu,\lambda}(\tau)$. The number of individuals $n$ will eventually be summed over.
@@
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other $n-1$ individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N$ individuals at the time of observation. For an illustration, see Figure 4.
@@
In Appendix .1 we show how this expression evaluates to $2\lambda\frac{p_{1}^{2}}{1-p_{0}}=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}$ and a normalisation constant of one. We find the same result for a final population size $N\geq 2$. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $\tau$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Running time forwards again and conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
@@ IX Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time, with a particular focus on how branch lengths change when the population is sampled at different sampling probabilities. We look at two distinct statistical scenarios: sampling at a particular time since the birth-death process started and sampling when the population has a given size $N$.
@@
For the first scenario, we generalize the results of Paradis Paradis (2016) to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. A second consequence of this behaviour under sampling concerns the number of generations along interior branches: even at low sampling probabilities, the number of generations along interior branches can be small — depending on the ratio of death and birth rates, but asymptotically independently of the sampling. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than accumulating continuously over time. The consequences and applications of this result will be explored in a future publication.
@@
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (like a branch of a given length), relative to the probability of a population with a certain size. The combination of going backwards and forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches.
@@
The interior branch distributions of the asymptotic cases of both scenarios (large times and large population sizes) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size N, even when N is not large, as the latter distribution is independent of $N$ Stadler and Steel (2012).
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant SFB1310/2 - 325931972. Many thanks to Tibor Antal, Samuel Johnston, Joachim Krug for discussions, and to Tanja Stadler for comments throughout and especially for helpful advice on section VII. We thank the computing center of the University of Cologne RRZK for computing time on the DFG-funded CHEOPS cluster (grant number INST 216/512/1FUGG).

### v3

@@ VII Branch lengths conditioned on the population size at the time of observation
First we go backwards in time from the time of observation to the time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation has been characterized by Ignatieva, Hein and Jenkins Ignatieva et al. (2020). It is described by that of a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{n,N}^{\mu,\lambda}(\tau)$. The number of individuals $n$ will eventually be summed over.
@@
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other $n-1$ individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N=2$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N=2$ individuals at the time of observation. For an illustration, see Figure 4.
@@
In Appendix .1 we show how this expression evaluates to $2\lambda\frac{p_{1}^{2}}{1-p_{0}}=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}$ and a normalisation constant of one. We find the same result for a final population size $N\geq 2$, see Appendix .1. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $\tau$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Running time forwards again while conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
@@ VIII Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time. We look at two distinct statistical scenarios: sampling at a given moment in time since the birth-death process started and sampling when to population has a given size $N$. A particular focus is on how branch lengths are affected when the population is sampled at different sampling probabilities and only a random subset of individuals can be used to reconstruct the tree.
@@
For the first scenario, we generalize the approach of Paradis Paradis (2016) of observing the population at a given time to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic distribution. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. An interesting consequence of sampling at a low probability concerns the number of generations (birth events which have not lead to additional sampled clades) along interior branches: even at low sampling probabilities, the number of generations along interior branches can be small due to the short branch lengths. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than by a continuous accumulation over time. The consequences and applications of this result will be explored in a future publication.
@@
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (such as a branch of a given length), relative to the probability of a population with a certain size. The combination of first going backwards and then forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). The ”going backwards in time”-aspect of this approach is shared with coalescent approaches, like the Kingman-coalescent Kingman (1982). However, the statistics we compute follows that of the original birth-death process which generated the tree. It thus includes the effects of lineages that died out (reflected for instance in the number of generations along a branch), and the distinct behaviour of pendant and interior branches under sampling, which are generally missing in coalescent approaches. Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches.
@@
The interior branch distributions of the asymptotic cases of both scenarios (large times for scenario i) and large population sizes for ii) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size $N$, even when $N$ is not large, as the latter distribution is independent of $N$ Stadler and Steel (2012).
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant SFB1310/2 - 325931972. Many thanks to Tibor Antal, Samuel Johnston, Joachim Krug, Thomas Wiehe for discussions, and to Tanja Stadler for comments throughout and especially for helpful advice on section VI. We thank the computing center of the University of Cologne RRZK for computing time on the DFG-funded CHEOPS cluster (grant number INST 216/512/1FUGG).
Appendix: Branch length distributions conditioned on the population size at the time of observation
The probability of a birth-death process with birth rate $\lambda$ and death rate $\mu$ taking a population of size $n>0$ to size $N\geq 0$ over time $t$ is given by (equation 8.47 in Bailey [1991])
$p_{N,n}^{\lambda,\mu}(t)$ $=\sum_{j=0}^{\min(n,N)}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}\alpha^{n-j}\beta^{N-j}(1-\alpha-\beta)^{j}$ (25)
with shorthand
$\alpha=\alpha(t,\lambda,\mu)$ $=\mu\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-1}{\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-\mu}$ (26)
$\beta=\beta(t,\lambda,\mu)$ $=\lambda\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-1}{\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)t}-\mu}\ .$ (27)
This probability can be derived by considering $n$ clades starting out with a single individual at a particular time and evolving according to (4), and computing the probability that the sum of all clade sizes a time $t$ later is $N$. For $N=0$ equation (25) simplifies to $p_{0,n}^{\lambda,\mu}(t)=\alpha^{n}=p_{0}(\tau)^{n}$.
The conditional probability (25) can also be used to describe that statistics of the population size $n$ some time $\tau$ in the past, conditioned on a population size $N$ at the time of observation. Ignatieva, Hein, and Jenkins Ignatieva et al. [2020] show that the statistics of $n$ is described by a birth-death process running for time $\tau$, starting with $N$ and with the rates of birth and death exchanged. This result uses a flat prior on the (unknown) age $T$ of the birth-death process. The same flat prior is also used explicitly in the derivation of the pendant branch length distribution conditioned on the final population size by Stadler and collaborators Mooers et al. [2012], Stadler and Steel [2012].
An important observation that simplifies the subsequent calculations is that the ratio of conditional probabilities going forwards and backwards in time simplifies to
$\frac{p_{n,N}^{\mu,\lambda}(\tau)}{p_{N,n}^{\lambda,\mu}(\tau)}=N/n\ .$ (28)
for all times and rates. This can be derived using the symmetry in the exchange of birth and death rates $\alpha(t,\mu,\lambda)=\beta(t,\lambda,\mu)$. Ignoring the binomial coefficients in equation (25), the remaining terms are therefore invariant to swapping $\lambda\leftrightarrow\mu$ and $N\leftrightarrow n$. For the binomial coefficients we get (${{N}\choose{j}}{{N+n-j-1}\choose{N-1}}=\frac{N}{n}{{n}\choose{j}}{{n+N-j-1}\choose{n-1}}$. Since the upper limit of the sum in equation (25) is invariant under exchange of $n$ and $N$, we get the result (28).
.1 Pendant branches
We now solve the toy problem of Section VII (pendant branch length distribution conditioned on a final population of size $N=2$) in three different ways of increasing generalizability. The last one generalizes directly to the result for any value of the final population size $N$ and to the goal of this appendix, the distribution of interior branch lengths conditioned on population size.
The first approach takes the sum (24), which with (28) becomes the straightforward geometric sum
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{n,2}^{\mu,\lambda}(\tau)}{p_{2,n}^{\lambda,\mu}(\tau)}=2\lambda p_{1}(\tau)^{2}\sum_{n=1}^{\infty}p_{0}(\tau)^{n-1}=2\lambda\frac{p_{1}(\tau)^{2}}{1-p_{0}(\tau)}\ .$ (29)
In the second approach, we look at all individuals just after the cell division event at time $\tau$ that establishes the pendant branch we are considering. Figure 5A sketches the situation: There are $n+1$ individuals, one of which has just established the pendant branch. This individual will have exactly one extant offspring at the time of observation. The other $n$ individuals must also have exactly one extant offspring in total, leading to a population size of $N=2$ at the time of observation. Additionally, the particular individual that was also born at time $\tau$ (but does not establish the branch we consider) must have at least one extant offspring. (For the case $N=2$, the branches starting with the birth event at time $\tau$ are equivalent, but these statements generalize to $N>2$). The probability of the $n$ individuals having one extant offspring at the time of observation, with one given individual not going extinct is $P(1|n)-p_{0}P(1|n-1)$ with shorthand $P(N|n)\equiv p_{N,n}^{\lambda,\mu}(\tau)$, $p_{0}=p_{0}(\tau)$. (Similarly, we do not note all dependencies of $\tau$ in the following equations, i.e. $p_{1}=p_{1}(\tau)$.) In this approach, we rewrite the relative fraction of birth events leading to a pendant branch at time $\tau$ (24) as
$\sum_{n=1}^{\infty}\lambda np_{1}(P(1|n)-p_{0}P(1|n-1))\frac{p_{n,2}^{\mu,\lambda}}{p_{2,n}^{\lambda,\mu}}=2\lambda p_{1}\sum_{n=1}^{\infty}(P(1|n)-p_{0}P(1|n-1))=2\lambda p_{1}(1-p_{0})\sum_{n=1}^{\infty}P(1|n)\ ,$ (30)
which gives the same result as above since $P(1|n)=np_{1}p_{0}^{n-1}$ and $\sum_{n=1}^{\infty}np_{1}p_{0}^{n-1}=p_{1}\partial_{p_{0}}\frac{p_{0}}{1-p_{0}}=\frac{p_{1}}{(1-p_{0})^{2}}$.
ARefer to caption               BRefer to caption
Figure 5: Formation of branches given population size $N$ (A) Given a population size $N$ at the time of observation and $n$ individuals at time $\tau$ in the past, a pendant branch (red) forms from a birth event where one offspring has exactly one and the other offspring has at least one surviving descendant. The number of descendants from the latter and the other $n-1$ individuals at time $\tau$ must equal $N-1$. (B) The same situation, but for an interior branch: At time $\tau_{s}$ one of the $n$ individuals undergoes a birth event. To form an interior branch (red), one offspring has to form a lineage until time $\tau_{e}$ where all additional lineages arising die out before the population is observed. At time $\tau_{e}$, the individual undergoes a birth event with both offspring forming observed clades with a total of $m\geq 2$ individuals. The second offspring from the birth event at $\tau_{s}$ cannot die out but forms a clade of size $r\geq 1$. The remaining $N-m-r$ observed individuals have to be offspring of the $n-1$ individuals that did not divide at time $\tau_{s}$.
The third approach uses a generating function to compute $P(1|n)$. The probability that a given clade has size $n$ at a time $\tau$ after it originated with a single individual is given by (2)-(4) and can be expressed as
$p_{n}(\tau)$ $=p_{n,1}^{\lambda,\mu}(\tau)=(1-\alpha)(1-\beta)\beta^{n-1}=p_{1}(\tau)\left(\frac{\lambda}{\mu}p_{0}(\tau)\right)^{n-1}$ $n$ $>0$
$p_{0}(\tau)$ $=p_{0,1}^{\lambda,\mu}(\tau)=\alpha$ $n$ $=0\ .$ (31)
The corresponding generating function is
$g(x)=\sum_{k=0}^{\infty}p_{k}(\tau)x^{k}=\alpha+(1-\alpha)(1-\beta)\sum_{k=1}^{\infty}\beta^{k-1}x^{k}\ .$ (32)
$P(1|n)$ is the coefficient of $x$ to the first power in the $n$th power of the generating function $g(x)$, which is $P(1|n)=\partial_{x}|_{x=0}g(x)^{n}=np_{1}p_{0}^{n-1}$ as in the second approach. However, since the geometric sum over $g(x)^{n}$ converges absolutely ($|\alpha|<1$ and $|\beta|<1$ for $\mu<\lambda$ and $\tau>0$), the summation over $n$ and differentiation commute, so we can perform the sum over $n$ first
$\sum_{n=1}^{\infty}P(1|n)=\sum_{n=1}^{\infty}\partial_{x}|_{x=0}g(x)^{n}=\partial_{x}|_{x=0}\sum_{n=1}^{\infty}g(x)^{n}=\partial_{x}|_{x=0}\frac{g(x)}{1-g(x)}\ .$ (33)
Straightforward algebra gives
$\frac{g(x)}{1-g(x)}=\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\ .$ (34)
Since $\frac{x}{1-x}=-1+\frac{1}{1-x}$, $\frac{x}{1-x}$ and $\frac{1}{1-x}$ have the same derivatives giving
$\sum_{n=1}^{\infty}P(1|n)=\partial_{x}|_{x=0}\left[\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\right]=\frac{1-\beta}{1-\alpha}=p_{1}/(1-p_{0})^{2}$ (35)
as before.
The third approach based on the generating function generalizes immediately to any given final population size. With (28) the relative fraction of birth events leading to a pendant branch is instead of (30)
$N\lambda p_{1}(1-p_{0})\sum_{n=1}^{\infty}P(N-1|n)$ (36)
and
$\sum_{n=1}^{\infty}P(N-1|n)=\frac{1}{(N-1)!}\partial_{x}^{N-1}|_{x=0}\left[\frac{\alpha}{1-\alpha}\frac{1}{1-x}+\frac{1-\alpha-\beta}{1-\alpha}\frac{x}{1-x}\right]=p_{1}/(1-p_{0})^{2}\ ,$ (37)
independently of the final population size $N$. The resulting normalization factor turns out to be $2/N$, which means that the distribution of pendant branch lengths conditioned on the population size $N$ is independent of $N$.
Superficially, the result (29) looks different from the distribution of pendant branch lengths derived in Stadler and Steel Stadler and Steel [2012] and Mooers et al. Mooers et al. [2012], which in the present notation is $2\lambda p_{1}(\tau)\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)$. The results agree, as they must, since $\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)=(1-\beta)=\frac{p_{1}(\tau)}{1-p_{0}(\tau)}$ (compare to (.1)). Furthermore, the result coincides with the distribution of pendant branches conditioned on large tree ages $T$ (10) derived in Section III, as $p_{1}(\tau)/(1-p_{0}(\tau))=\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}(1-p_{0}(\tau))$.
.2 Interior branches
The approach based on generating functions can also be applied straightforwardly to the statistics of interior branches conditioned on the final population size $N$. Figure 5B shows how a population of size $n$ at time $\tau_{s}$ gives rise to an interior branch. The branch we focus on must have $m\geq 2$ extant offspring at the time of observation (with only one offspring it would be a pendant branch), the second individual born at time $\tau_{s}$ must have $r\geq 1$ offspring, and the remaining $n-1$ individuals have in total $N-r-m$ offspring. The latter condition ensures that the final population size equals $N$. The probability for such a trajectory starting from $n+1$ individuals at time $\tau_{s}$ and ending with $N$ at the time of observation is
$p_{\mbox{mr}}(\tau_{s})=P(N|n+1)-\left(p_{0}P(N|n)+p_{1}P(N-1|n)\right)-p_{0}P(N|n)+\left(p_{0}p_{0}P(N|n-1)+p_{1}p_{0}P(N-1|n-1)\right)\ ,$ (38)
where the suppressed time is $\tau_{s}$. The first term gives the probability for ending up with $N$ individuals when starting with $n+1$, the next term in brackets excludes the two cases $m=0$ and $m=1$, and the term following that excludes the case $r=0$. The last term in brackets corrects overcounting the case when both $m<2$ and $r=0$.
The next step looks at the clade establishing the interior branch at time $\tau_{s}$. Given that this clade has size $m\geq 2$ at the time of observation, what is the probability density (per small time period) of a birth event at time $\tau_{e}$, while the branch has no offspring with extant descendants prior to $\tau_{e}$ but is terminated at $\tau_{e}$, because both offspring establish extant lineages. This is
$\frac{p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}}{1-p_{0}(\tau_{s})-p_{1}(\tau_{s})}\ ,$ (39)
where the denominator arises from conditioning on two extant or offspring $m\geq 2$.
Putting these results together, the probability density (up to a normalizing factor) at which interior branches start at time $\tau_{s}$ and end at time $\tau_{e}$ conditioned on a population size $N$ at the time of observation is
$\frac{p_{1}^{\tau_{e}}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}}{1-p_{0}(\tau_{s})-p_{1}(\tau_{s})}\sum_{n=1}^{\infty}\lambda n\frac{p_{n,N}^{\mu,\lambda}(\tau_{s})}{p_{N,n}^{\lambda,\mu}(\tau_{s})}p_{\mbox{mr}(\tau_{s})}\ .$ (40)
Although this expression looks cumbersome, it can be easily evaluated using (28) and (37). Using $P(N|0)=0$ and $\sum_{n=1}^{\infty}P(N|n+1)=\sum_{n=1}^{\infty}P(N|n)-P(N|1)$, the sum reduces to
$\sum_{n=1}^{\infty}p_{\mbox{mr}}(\tau_{s})$ $=(1-2p_{0}(\tau_{s})-p_{1}(\tau_{s})+p_{0}(\tau_{s})p_{0}(\tau_{s})+p_{1}(\tau_{s})p_{0}(\tau_{s}))\frac{p_{1}(\tau_{s})}{(1-p_{0}(\tau_{s}))^{2}}-p_{N}(\tau_{s})$
$=(1-p_{0}(\tau_{s})-p_{1}(\tau_{s}))\frac{p_{1}(\tau_{s})}{1-p_{0}(\tau_{s})}-p_{N}(\tau_{s})\ .$ (41)
Thus, using equation (.1) to replace $p_{N}(\tau_{s})$, the probability density of an internal branch is proportional to
$\lambda^{2}Np_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\left(1-p_{0}(\tau_{e})\right)^{2}\left[\frac{p_{1}(\tau_{s})}{1-p_{0}(\tau_{s})}-\frac{p_{1}(\tau_{s})\left(\frac{\lambda}{\mu}p_{0}(\tau_{s})\right)^{N-1}}{1-p_{0}(\tau_{s})-p_{1}(\tau_{s})}\right]\ .$ (42)
The second term in the square brackets vanishes quickly with increasing population sizes. When neglecting this term, the normalizing factor again turns out to be $2/N$, removing the dependency on $N$ entirely. Hence, up to a term of order $\mathcal{O}(p_{0}(\tau_{s})^{N-1})$, the distribution of interior branch lengths conditioned on the population size $N$ coincides with probability density function (13), the asymptotic distribution in trees conditioned on large $T$ as $p_{1}(\tau_{s})/(1-p_{0}(\tau_{s}))=\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}(1-p_{0}(\tau_{s}))$.
.3 Finite sampling
The results conditioned on a fixed final population size can be generalized to the finite sampling probability $\rho$ discussed in Section VI. Again, $p_{0}(\tau)$, $p_{1}(\tau)$ and $p_{1}^{(\tau_{e})}(\tau)$ are replaced by $p_{0}^{(\rho)}(\tau)$, $p_{1}^{(\rho)}(\tau)$ and $p_{1}^{(\rho)(\tau_{e})}(\tau)$ (equations (15)-(17)). However, we do not know the population size $N$, only the number of samples.
In the case of pendant branches, this means we do not know the exact number of extant descendants left by the $n$ individuals after the birth event at time $\tau$ who do not form the pendant branch (in the case of complete sampling they need to be $N-1$). However, since equation (37) is independent of $N$, summing over all possible numbers of descendants leads to the same factor, and the density function of pendant branches is thus (up to normalization) given by equation (29) with $p_{0}^{(\rho)}(\tau)$ and $p_{1}^{(\rho)}(\tau)$, coinciding with the result for trees at a given large age $T$ of the birth-death process in the case of finite sampling (18).
For the interior branch distribution, introducing finite sampling with $p_{0}^{(\rho)}(\tau)$, $p_{1}^{(\rho)}(\tau)$ and $p_{1}^{(\rho)(\tau_{e})}(\tau)$ covers the asymptotic (and N-independent) case of large $N$, which again coincides with the probability density functions conditioned on large $T$ (19)-(20). The correction term for small $N$, however, is more challenging to generalize to finite sampling since $N$ is not known. In principle, it can be treated by summing over $N$ conditioned on a fixed number of samples and then following through with the steps in Appendix .2.